OpenStudy (anonymous):

is it true in an exponential problem if your initial starting number is 2 that you have to double it to get the next number?

OpenStudy (bahrom7893):

true

OpenStudy (anonymous):

=)

OpenStudy (bahrom7893):

2^n is the exponential, 2^1 = 2 - starting 2^2 = 2*2 double the starting 2^3 = (2*2)*2 double again, etc..

OpenStudy (anonymous):

ahhhhhhh i got this =) your really good at math

OpenStudy (bahrom7893):

i'm a math major and i suck at most of upper level math classes

OpenStudy (anonymous):

nice . . . . really? it doesnt seem like it . . .

OpenStudy (bahrom7893):

well u don't get advanced questions asked here.. usually people can't answer the questions that I ask. Satellite and a couple of others can help with half of my hw, but not all of it haha

OpenStudy (anonymous):

ohhh hahahahahaha